Suppose that price of AAA stock at the end of next year depends on the state of the economy. The economy can have three states with the supplied probabilities and prices. What is the standard deviation of AAA’s returns if you purchase AAA today for $100? Also assume that AAA will pay a dividend of $10 at the end of the year. Enter your answer as a percent without the % sign. Round your final answer to 2 decimals.

Expansion

Stable

Recession

Probability

.25

.5

.25

Price in one year

$130

$100

$80
Answer #1

Date: JU Calculation of Standard Deuiation of Returns 1st we will Calculate Return Return - Dividend & Share Price at - PurcNow Praledeility Return line I Return (ime) (x) dx= x-1 (=20-50 (dac² CP .25 27.50 Expaniam Stable Recessian 40 lo 189.0625 .